解析的整数論とその周辺

研究集会

京都大学数理解析研究所の共同研究事業の一つとして、下記のように研究集会を催しますので、ご案内申し上げます。

研究代表者松本耕二 (名古屋大学 · 多元数理)

記日時： 1999 年 11 月 29 日 (月) 10:30 ∼

12 月 3 日 (金) 16:30 場所:京都大学数理解析研究所 4 階 420 号室

京都市左京区北白川追分町市バス農学部前または北白川下車

プログラム

11 月 29 日 (月)

10:30 ∼ 10:35 Opening

10:35 ∼ 11:15 江上繁樹 (富山大 · 工) Some identities associated to weighted mean values of Dirichlet series

11:20 ∼ 12:00 永田誠 (京大 · 数理研) G-関数についての最近の結果 (Recent results on G-functions)

14:00 ∼ 14:40 天羽雅昭 (群馬大 · 工) 桂田昌紀 (慶応大 · 経済) On the values of certain q-hypergeometric series

14:45 ∼ 15:15 神谷諭一 (名大 · 多元数理) Certain mean values and non-vanishing of automorphic L-functions with large level

15:20 ∼ 15:50 石川秀明 (新潟大 · 自然科学) On a sum related to a multiple L-function

16:10 ∼ 16:40 T. P. Peneva (筑波大 · 数学) On the exceptional set in Goldbach’s problem

16:45 ∼ 17:35 田村純一 Certain tilings and words of higher dimension, and their non-periodicity

11 月 30 日 (火)

9:30 ∼ 10:20 北岡良之 (名大 · 多元数理)

実 2 次体の単数の分布について (On the distribution of units of real quadratic fields)

10:25 ∼ 11:15 Chaohua Jia (Academia Sinica) The application of the combination of circle method and sieve method

11:20 ∼ 12:00 安富真一 (鈴鹿工業高専)

Dynamical representations of substituted Sturmian sequences

14:00 ∼ 14:30 若林功 (成蹊大 · 工)

On a family of cubic Thue inequalities

14:35 ∼ 15:05 三河寛 (筑波大 · 数学)

等差数列における最小素数について (On the least prime in an arithmetic progression)

15:10 ∼ 15:40 後藤和雄 (鳥取大 · 教育)

大久保幸夫 (鹿児島経済大 · 経済) Discrepancy of some special sequences

16:00 ∼ 16:30 吉元昌己 (九大 · 数理科学)

Zeta-regularization and Kronecker’s limit formulas

16:35 ∼ 17:05 市原由美子 (名大 · 多元数理)

The evaluation of the sum over arithmetic progressions for the coefficients of the Rankin-Selberg series

17:10 ∼ 17:40 見正秀彦 (名大 · 多元数理)

イデアル類群の指標を持つ L 関数の universality (Universality of L-functions with characters of ideal class groups)

12 月 1 日 (水)

9:30 ∼ 10:20 J. Brüdern (Universität Stuttgart) Multiplicative sequences, binary additive problems, and the Bateman-Horn conjecture

10:25 ∼ 11:15 西岡久美子 (慶応大 · 経済)

A vanishing theorem for Mahler functions

11:20 ∼ 12:10 Wenpeng Zhang (Northwest University) On the 2k-th power mean of Dirichlet L-functions with the weight of general Kloosterman sums

12 月 2 日 (木)

9:30 ∼ 10:20 本橋洋一 (日本大 · 理工)

From Hecke to Riemann via spherical harmonics

10:25 ∼ 11:15 A. Sankaranarayanan (Tata Institute of Fundamental Research) On the estimation of twisted exponential sum related to certain arithmetic functions

11:20 ∼ 12:00 小松尚夫 (三重大 · 教育)

On inhomogeneous Diophantine approximation and the Borweins’ algorithm

14:00 ∼ 14:40 岡崎龍太郎 (同志社大 · 工) On a gap principle for exponential Diophantine equation

14:45 ∼ 15:15 村林直樹 (山形大 · 理)

梅垣敦紀 (早稲田大 · 理工) Determination of all Q-rational CM-points in the moduli space of principally polarized abelian surfaces

15:20 ∼ 15:40 鹿野健 (山形大 · 教育)

Properties of Iwaniec’s q-functions

16:00 ∼ 16:30 佐野友紀 (津田塾大 · 情報数理) 伊藤俊次 (津田塾大 · 情報数理) On periodic β-expansions of Pisot numbers and Rauzy fractals

16:35 ∼ 17:05 利光剛 (慶応大 · 理工)

Algebraic solutions of a certain type of functional equations

17:10 ∼ 17:40 熊谷博 (近畿大 · 九州工)

Rapidly convergent series expressions for the zeta- and L-values

12 月 3 日 (金)

9:30 ∼ 10:20 S. Louboutin (Université de Caen) Bounds for |L(1, χ)| and Res_{s=1}(ζ_{K})

10:25 ∼ 11:15 川田浩一 (岩手大 · 教育)

Topics in Waring’s problem for fourth powers II 11:20 ∼ 12:00 寺井伸浩 (足利工大 · 工)

On the Diophantine equations on binomial coefficients

14:00 ∼ 14:40 秋山茂樹 (新潟大 · 理)

On the tiling generated by quadratic canonical number system

14:45 ∼ 15:15 原瀬巍 (鳥取大 · 教育)

A remark on the transcendency and the rationality of formal power series over finite fields

15:20 ∼ 15:50 日比野剛士 (早稲田大 · 理工総研) 菅真紀子 (お茶の水女子大 · 人間文化) θ-congruent numbers and modular parametrizations

15:55 ∼ 16:25 松本耕二 (名大 · 多元数理) Asymptotic expansions of double zeta-functions of Barnes, of Shintani, and Eisenstein series